Circular No. 8032

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

CBAT@CFA.HARVARD.EDU (science)

URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2002 X3 AND C/2002 X4 (SOHO) Further to IAUC 8026, D. Hammer reports measurements for two

Kreutz sungrazing comets found by X.-M. Zhou on SOHO website C2 coronagraph images that were taken around the time of the total solar eclipse. The reduced astrometry (discovery positions below) and orbital elements by B. G. Marsden appear on MPEC 2002-X64.

Comet  2002 UT  α_{2000}  δ_{2000}
C/2002 X3 Dec. 4.171  16^{h}43.^{m}1  −24^{o}03′
C/2002 X4  4.254  16 43.6  −24 03

COMET 2002 X5

S. Nakano, Sumoto, Japan, reports the visual discovery of a comet by Tetuo Kudo. The following observations have been forwarded by Nakano:

2002 UT  α_{2000}  δ_{2000}  m_{1} Observer
Dec. 13.833  15^{h}44^{m}  +45^{o}30′  9.5 Kudo
14.783  15 52  +44 57  9.5  _{”}
14.86146 15 53 07.^{s}63 +44 52 43′′.3 9.6 Kadota
14.86205 15 53 07.94 +44 52 42.3  _{”}
14.86278 15 53 08.31 +44 52 40.9  _{”}
14.86302 15 53 08.48 +44 52 39.9 7.5 Kushida
14.86338 15 53 08.59 +44 52 40.1  Kadota
14.86410 15 53 08.93 +44 52 38.3  _{”}
14.86458 15 53 09.18 +44 52 36.0  Kushida
14.86473 15 53 09.25 +44 52 37.3  Kadota
14.86597 15 53 09.87 +44 52 35.0  Kushida
14.86806 15 53 10.60 +44 52 31.1  _{”}

T. Kudo (Nishi Goshi-machi, Kikuchi-gun, Kumamoto-ken, Japan). Nikon 20×120 binoculars. Coma diameter 2′; central condensation visible, but no tail detected.

K. Kadota (Ageo, Saitama, Japan). 0.25-m f/5.0 reflector + CCD. Coma

diameter 5′.5; tail at least 18′ long in p.a. 331^{o}. R. and Y. Kushida (Yatsugatake South Base Observatory). 0.40-m f/10

Schmidt-Cassegrain reflector + CCD. Coma diameter 4′; tail at least 15′ in p.a. 330^{o}.

2002 December 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green